OpenStudy (anonymous):

What is the ‘d’ value for the following sequence: 23, 27, 31, 35…?

OpenStudy (mathstudent55):

Subtract any term from the very next term.

OpenStudy (anonymous):

ok so like 31-28=3

OpenStudy (mathstudent55):

Yes, but be careful, because there is no 28 in your sequence.

OpenStudy (mathstudent55):

27 - 23 = ? 31 - 27 = ? 35 - 31 = ? All the above differences are equal and are the "d" value you are looking for.